# Integrable measure equivalence rigidity of right-angled Artin groups via quasi-isometry

### v1

@@ 3.1 Blow-up buildings
A metric simplicial graph $\beta$ is a branched line if there exists $C>0$ such that $\beta$ is obtained from $\mathbb{R}$ (equipped with its simplicial structure given by subdividing at integer points) by gluing, at every integer $n\in\mathbb{Z}$, at most $C$ edges of length $1$, denoted $e_{n,1},\dots,e_{n,k}$, gluing the origin of each $e_{n,i}$ at $n$. Valence one vertices of $\beta$ are called the tips of $\beta$, and their set is denoted by $t(\beta)$. The copy of $\mathbb{R}$ is called the core of $\beta$. A branched flat $F$ is a product of finitely many branched lines $\beta_{1},\dots,\beta_{k}$. A tip of $F$ is a tuple $(t_{1},\dots,t_{k})$, where each $t_{i}$ is a tip of $\beta_{i}$; we denote their set by $t(F)$. The core of $F$ is the product of the cores of the $\beta_{i}$.
@@ 3.2 Properties of blow-up buildings
We now define a projection map $\pi:Y\to\mathbb{B}_{\Gamma}$. Note that for each standard line $\ell\subset G$, we can define a map $\pi:\beta_{\ell}\to\mathbb{B}_{\Gamma}$ by sending the core of $\beta_{\ell}$ to the rank 1 vertex in $\mathbb{B}_{\Gamma}$ associated with $\ell$, sending each vertex in $t(\beta_{\ell})$ to the associated rank 00 vertex in $\mathbb{B}_{\Gamma}$, and extending linearly. More generally, let $F=\prod_{\mathsf{v}\in V(\Delta(F))}\ell_{\mathsf{v}}$ be a standard flat, where each $\ell_{\mathsf{v}}\subset F$ is a standard $\mathsf{v}$-line, and let $\beta_{F}=\prod_{\mathsf{v}\in V(\Delta(F))}\beta_{\mathsf{v}}$ be the associated branched flat. We define $\pi:\beta_{F}\to\mathbb{B}_{\Gamma}$ as follows. Every vertex $x\in\beta_{F}$ lies in the core of a unique subcomplex of the form $\beta_{F^{\prime}}$, with $F^{\prime}\subseteq F$, and we let $\pi(x)$ be the vertex of $\mathbb{B}_{\Gamma}$ associated to $F^{\prime}$. These maps defined on each $\beta_{F}$ are compatible with the gluing pattern, hence induce a map $\pi:Y\to\mathbb{B}_{\Gamma}$. Note that the restriction of $\pi:Y\to\mathbb{B}_{\Gamma}$ to each cube is either an isometry or collapses the cube to a cube of smaller dimension (by collapsing some of the interval factors). We say that a vertex $y\in Y$ has rank $k$ if $\pi(y)$ has rank $k$. We record the following properties of $Y$.
@@
    There exists a unique injective map $f:G\to Y$ whose restriction to any standard line $\ell$ coincides with $f_{\ell}$. This map $f$ sends the vertex set of each standard flat of $G$ bijectively to the tips of a standard branched flat. The image of $f$ is exactly the set of 00-dimensional standard flats in $Y$.
@@
Let $(g_{\ell})_{\ell}$ be a uniformly locally finite blow-up datum, let $(f_{\ell})_{\ell}$ be an adapted family of bijections, and let $Y$ be the associated blow-up building. Then $G$ and $Y$ are quasi-isometric.
@@
Let $\alpha:H\curvearrowright\mathbb{B}_{\Gamma}$ be an action by cubical automorphisms. Let $(g_{\ell})_{\ell}$ be a blow-up datum, let $(f_{\ell})_{\ell}$ be an adapted family of bijections, and let $Y$ be the associated blow-up building. If $\alpha$ is compatible with $(g_{\ell}),(f_{\ell})$, then there exists an action $\alpha^{\prime}:H\curvearrowright Y$ by cellular isometries such that the map $\pi:Y\to\mathbb{B}_{\Gamma}$ is $(\alpha^{\prime},\alpha)$-equivariant.
@@ 3.3 Proof of the quasi-isometry criterion
Let $\{\mathsf{v}_{1},\dots,\mathsf{v}_{n}\}$ be a (finite) set of representatives of the orbits of vertices for the action $H\curvearrowright\Gamma^{e}$. By assumption, the action $\alpha_{\mathsf{v}_{i}}$ is conjugate to an action on $\mathbb{Z}$ by uniform quasi-isometries. Therefore, by [HK18, Proposition 6.3], it is semi-conjugate to an action on $\mathbb{Z}$ by isometries. More precisely, there exist an isometric action $\gamma_{\mathsf{v}_{i}}:H_{\mathsf{v}_{i}}\curvearrowright\mathbb{Z}$, an $(\alpha_{\mathsf{v}_{i}},\gamma_{\mathsf{v}_{i}})$-equivariant surjection $g_{i}:Z_{\mathsf{v}_{i}}\to\mathbb{Z}$, and $C_{i}>0$ satisfying $|g^{-1}_{i}(n)|\leq C_{i}$ for every $n\in\mathbb{Z}$. For every standard line $\ell$, we now define a map $g_{\ell}:\ell\to\mathbb{Z}$, following the construction in [HK18, Section 5.6]. First, if $\ell$ is a $\mathsf{v}_{i}$-line for some $i\in\{1,\dots,n\}$, we let $p_{i}:P_{\mathsf{v}_{i}}\to Z_{\mathsf{v}_{i}}$ be the projection map, and let $g_{\ell}=(g_{i}\circ p_{i})_{|\ell}$. In general $\ell$ is a $\mathsf{w}$-line for some $\mathsf{w}\in V\Gamma^{e}$. For every $\mathsf{w}\in V\Gamma^{e}$, we choose an element $h_{\mathsf{w}}\in H$ such that $h_{\mathsf{w}}\mathsf{w}=\mathsf{v}_{i}$ for some $i\in\{1,\dots,n\}$. Now if $\ell$ is a $\mathsf{w}$-line, then $h_{\mathsf{w}}(\ell)$ is a $\mathsf{v}_{i}$-line, and we let $g_{\ell}=g_{h_{\mathsf{w}}(\ell)}\circ h_{\mathsf{w}}$. As the maps $g_{i}$ are finite-to-one, the family $(g_{\ell})_{\ell}$ is a uniformly locally finite blow-up datum. Let $(\beta_{\mathsf{v}})_{\mathsf{v}\in V\Gamma^{e}}$ be the family of branched lines associated to the blow-up datum $(g_{\ell})_{\ell}$, and let $(f_{\ell}:\ell\to t(\beta_{\mathsf{v}}))_{\ell}$ be an adapted family of bijections. For every $i\in\{1,\dots,n\}$, the equivariance property of the map $g_{i}$ gives an isometric action $\alpha^{\prime}_{\mathsf{v}_{i}}:H_{\mathsf{v}_{i}}\curvearrowright\beta_{\mathsf{v}_{i}}$. More generally, for every $\mathsf{w}\in V\Gamma^{e}$, there is an isometric action $\gamma_{\mathsf{w}}:H_{\mathsf{w}}\curvearrowright\mathbb{Z}$ given by $\gamma_{\mathsf{w}}(h)(z)=\gamma_{\mathsf{v}_{i}}(h_{\mathsf{w}}hh_{\mathsf{w}}^{-1})(z)$ (where $i$ is such that $h_{\mathsf{w}}\mathsf{w}=\mathsf{v}_{i}$). Then for every $\mathsf{w}$-line $\ell$, the map $g_{\ell}:\ell\to\mathbb{Z}$ constructed above is $(\alpha_{\mathsf{w},\ell},\gamma_{\mathsf{w}})$-equivariant. So it yields an isometric action $\alpha^{\prime}_{\mathsf{w}}:H_{\mathsf{w}}\curvearrowright\beta_{\mathsf{w}}$, such that the map $f_{\ell}:\ell\to t(\beta_{\mathsf{w}})$ is $(\alpha_{\mathsf{w},\ell},\alpha^{\prime}_{\mathsf{w}})$-equivariant. One also checks that the second compatibility condition in Definition 3.7 is satisfied. Let $Y$ be the blow-up building associated to $(g_{\ell})_{\ell},(f_{\ell})_{\ell}$. By Lemma 3.6, the space $Y$ is quasi-isometric to $G$. By Lemma 3.8, there is an action $\alpha^{\prime}:H\curvearrowright Y$ by cellular isometries, such that the map $\pi:Y\to\mathbb{B}_{\Gamma}$ is $(\alpha^{\prime},\alpha)$-equivariant. We now prove that the action $\alpha^{\prime}$ is proper and cocompact. This will imply that $H$ is finitely generated and quasi-isometric to $G$, as desired. By definition of $\pi$, every rank 00 vertex of $\mathbb{B}_{\Gamma}$ has a unique preimage under $\pi$. As the map $\pi:Y\to\mathbb{B}_{\Gamma}$ is $(\alpha^{\prime},\alpha)$-equivariant, it follows that the action $\alpha^{\prime}$ has finitely many orbits of rank 0 vertices. Note that every vertex $y$ of $Y$ of rank at least $1$ is adjacent to at least one vertex of lower rank (this follows by considering the standard flat containing $y$ and the associated standard branched flat). Thus there exists $C>0$ such that any vertex in $Y$ is at most distance $C$ from a rank 0 vertex. As $Y$ is uniformly locally finite, it follows that $\alpha^{\prime}$ has finitely many orbits of vertices. Using again that $Y$ is uniformly locally finite, this is enough to ensure that there are only finitely many orbits of cells, so $\alpha^{\prime}$ is cocompact. As $H$ acts cocompactly on a uniformly locally finite complex, to show that the action is proper, it suffices to show that the stabilizer of each vertex is finite. The case of rank 0 vertices follows from the first assumption of the theorem, the equivariance of $\pi$, and the fact that $\pi$ is a bijection between rank 00 vertices of $Y$ and $\mathbb{B}_{\Gamma}$. The equivariance of $\pi$ ensures that the $H$-action on $Y$ preserves the rank of vertices. Therefore, the stabilizer of each vertex $y\in Y$ of rank at least $1$ permutes the non-empty finite set of vertices of lower rank which are adjacent to $y$. By induction on the rank, we thus deduce that the stabilizer of every vertex is finite.
@@ 4.2 Reduction of couplings
In this section, we review work of Kida [Kid10, Kid11] and Bader–Furman–Sauer [BFS13]. Let $L$ be a Polish group, and let $\mathsf{G}$ be a countable subgroup of $L$. Then $L$ is equipped with an action of $\mathsf{G}\times\mathsf{G}$ by left-right multiplication, namely $(g_{1},g_{2})\cdot\ell=g_{1}\ell g_{2}^{-1}$. We say that the inclusion $\mathsf{G}\subseteq L$ is strongly ICC if the Dirac mass at $\mathrm{id}$ is the unique probability measure on $L$ which is invariant under the conjugation by every element of $\mathsf{G}$.
@@
Let $\mathsf{H}$ be a countable group that is measure equivalent to $\mathsf{G}$, and let $\Omega$ be a measure equivalence coupling between $\mathsf{G}$ and $\mathsf{H}$. Then there exist a homomorphism $\iota:\mathsf{H}\to L$ with finite kernel, and a measurable almost $(\mathsf{G}\times\mathsf{H})$-equivariant map $\theta:\Omega\to L$, i.e. for a.e. $\omega\in\Omega$, and any $(g,h)\in\mathsf{G}\times\mathsf{H}$, one has $\theta((g,h)\cdot\omega)=g\theta(\omega)\iota(h)^{-1}$.
@@
Then $\Phi$ is almost $(\hat{\mathsf{G}}\times\hat{\mathsf{G}})$-equivariant.
@@ 4.3 Restricting couplings to stabilizers
Let $L$ be a Polish group and $\mathsf{G}$ be a countable subgroup of $L$. Let $L^{\prime}\subseteq L$ be a Borel subgroup such that $\mathsf{G}\cdot L^{\prime}=L$. Let $\mathsf{H}$ be a countable group that is measure equivalent to $\mathsf{G}$, let $(\Omega,\mu)$ be a measure equivalence coupling between $\mathsf{G}$ and $\mathsf{H}$, and assume we are given a homomorphism $\iota:\mathsf{H}\to L$ and a measurable almost $(\mathsf{G}\times\mathsf{H})$-equivariant map $\theta:\Omega\to L$, where the $(\mathsf{G}\times\mathsf{H})$-action on $L$ is via $(g,h)\cdot\ell=g\ell\iota(h)^{-1}$. Then the groups $\mathsf{G}^{\prime}=\mathsf{G}\cap L^{\prime}$ and $\mathsf{H}^{\prime}=\iota^{-1}(L^{\prime})$ are measure equivalent. More precisely $\Omega^{\prime}=\theta^{-1}(L^{\prime})$ is a measure equivalence coupling between $\mathsf{G}^{\prime}$ and $\mathsf{H}^{\prime}$. In addition, for every subgroup $\mathsf{K}$ of either $\mathsf{G}$ or $\mathsf{H}$, every Borel fundamental domain for the action of $\mathsf{K}\cap\mathsf{G}^{\prime}$ (or $\mathsf{K}\cap\mathsf{H}^{\prime}$) on Ωsuperscriptnormal-$\Omega^{\prime}$ is contained in a Borel fundamental domain for the action of $\mathsf{K}$ on $\Omega$.
@@
Let $\mathsf{K}\subseteq\mathsf{H}$ be a subgroup, let $\mathsf{K}^{\prime}=\mathsf{K}\cap\mathsf{H}^{\prime}$, and let $Y^{\prime}$ be a Borel fundamental domain for the action of $\mathsf{K}^{\prime}$ on $\Omega^{\prime}$: this exists because the $\mathsf{H}$-action on $\Omega$ has one. We claim that for any $k\in\mathsf{K}\setminus\{1\}$, one has $\mu(kY^{\prime}\cap Y^{\prime})=0$. Indeed, for a.e. $x,y\in Y^{\prime}$, if $y=kx$ for some $k\in\mathsf{K}$, then $\theta(y)=\theta(x)\iota(k)^{-1}$. As $\theta(x),\theta(y)\in L^{\prime}$, we have $k\in\mathsf{K}^{\prime}$. The claim thus follows from the fact that $Y^{\prime}$ is a fundamental domain for the action of $\mathsf{K}^{\prime}$ on $\Omega^{\prime}$. The same argument applies to subgroups of $\mathsf{G}$. When $\mathsf{K}=\mathsf{H}$, the above claim implies in particular that every Borel fundamental domain for the action of $\mathsf{H}^{\prime}$ on $\Omega^{\prime}$ has finite measure. Likewise, any Borel fundamental domain for the action of $\mathsf{G}^{\prime}$ on $\Omega^{\prime}$ has finite measure.
@@
Let $K$ be a polyhedral complex with countably many cells. Let $L$ be a Polish group acting on $K$ through a measurable homomorphism $L\to\mathrm{Aut}(K)$. Let $\mathsf{G}$ be a countable subgroup of $L$, and assume that the actions of $\mathsf{G}$ and $L$ on $VK$ have the same orbits. Let $\mathsf{H}$ be a countable group that is measure equivalent to $\mathsf{G}$, let $(\Omega,\mu)$ be a measure equivalence coupling between $\mathsf{G}$ and $\mathsf{H}$, and assume we are given a homomorphism $\iota:\mathsf{H}\to L$ and a measurable almost $(\mathsf{G}\times\mathsf{H})$-equivariant map $\theta:\Omega\to L$. Then for every $v\in VK$, the groups $\mathsf{G}_{v}=\mathrm{Stab}_{\mathsf{G}}(v)$ and $\mathsf{H}_{v}=\iota^{-1}(\mathrm{Stab}_{L}(v))$ are measure equivalent. More precisely $\Omega_{v}=\theta^{-1}(\mathrm{Stab}_{L}(v))$ is a measure equivalence coupling between $\mathsf{G}_{v}$ and $\mathsf{H}_{v}$, and every Borel fundamental domain for the action of $\mathsf{G}_{v}$ (resp. $\mathsf{H}_{v}$) on $\Omega_{v}$ is contained in a Borel fundamental domain for the action of $\mathsf{G}$ (resp. $\mathsf{H}$) on $\Omega$.
@@
Given a countable set $\mathbb{D}$, we denote by $\mathrm{Bij}(\mathbb{D})$ the group of all bijections of $\mathbb{D}$; we equip it with the topology of pointwise convergence, for which it is a Polish group. Recall from the introduction that a group $\mathsf{H}$ has bounded torsion if there is a bound on the cardinality of a finite subgroup of $\mathsf{H}$.
@@
Let $\mathsf{H}$ be a countable group with bounded torsion that is measure equivalent to $\mathsf{G}$, and let $\Omega$ be a measure equivalence coupling between $\mathsf{G}$ and $\mathsf{H}$. Assume that we are given a homomorphism $\iota:\mathsf{H}\to L^{0}$ and a measurable almost $(\mathsf{G}\times\mathsf{H})$-equivariant map $\theta:\Omega\to L^{0}$.

### v2

@@ 3.1 Blow-up buildings
A metric simplicial graph $\beta$ is a branched line (see Figure 2 left) if there exists $C>0$ such that $\beta$ is obtained from $\mathbb{R}$ (equipped with its simplicial structure given by subdividing at integer points) by gluing, at every integer $n\in\mathbb{Z}$, at most $C$ edges and at least one edge of length $1$, denoted $e_{n,1},\dots,e_{n,k}$, gluing the origin of each $e_{n,i}$ at $n$.
Valence one vertices of $\beta$ are called the tips of $\beta$, and their set is denoted by $t(\beta)$. The copy of $\mathbb{R}$ is called the core of $\beta$.
A branched flat $F$ is a product of finitely many branched lines $\beta_{1},\dots,\beta_{k}$. A tip of $F$ is a tuple $(t_{1},\dots,t_{k})$, where each $t_{i}$ is a tip of $\beta_{i}$; we denote their set by $t(F)$. The core of $F$ is the product of the cores of the $\beta_{i}$.
Figure 2: A branched line and a projection map.
@@ 3.2 Properties of blow-up buildings
We now define a projection map $\pi:Y\to\mathbb{B}_{\Gamma}$. Note that for each standard line $\ell\subset G$, we can define a map $\pi:\beta_{\ell}\to\mathbb{B}_{\Gamma}$ by sending the core of $\beta_{\ell}$ to the rank 1 vertex in $\mathbb{B}_{\Gamma}$ associated with $\ell$, sending each vertex in $t(\beta_{\ell})$ to the associated rank 0 vertex in $\mathbb{B}_{\Gamma}$, and extending linearly (see Figure 2). More generally, let $F=\prod_{\mathsf{v}\in V(\Delta(F))}\ell_{\mathsf{v}}$ be a standard flat, where each $\ell_{\mathsf{v}}\subset F$ is a standard $\mathsf{v}$-line, and let $\beta_{F}=\prod_{\mathsf{v}\in V(\Delta(F))}\beta_{\mathsf{v}}$ be the associated branched flat. We define $\pi:\beta_{F}\to\mathbb{B}_{\Gamma}$ as follows. Every vertex $x\in\beta_{F}$ lies in the core of a unique subcomplex of the form $\beta_{F^{\prime}}$, with $F^{\prime}\subseteq F$, and we let $\pi(x)$ be the vertex of $\mathbb{B}_{\Gamma}$ associated to $F^{\prime}$. These maps defined on each $\beta_{F}$ are compatible with the gluing pattern, hence induce a map $\pi:Y\to\mathbb{B}_{\Gamma}$. Note that the restriction of $\pi:Y\to\mathbb{B}_{\Gamma}$ to each cube is either an isometry or collapses the cube to a cube of smaller dimension (by collapsing some of the interval factors). We say that a vertex $y\in Y$ has rank $k$ if $\pi(y)$ has rank $k$. We record the following properties of $Y$.
@@
    There exists a unique injective map $f:G\to Y$ whose restriction to any standard line $\ell$ coincides with $f_{\ell}$. This map $f$ sends the vertex set of each standard flat of $G$ bijectively to the tips of a standard branched flat. The image of $f$ is exactly the set of 0-dimensional standard flats in $Y$.
@@
Let $(g_{\ell})_{\ell}$ be a uniformly locally finite blow-up datum, let $(f_{\ell})_{\ell}$ be an adapted family of bijections, and let $Y$ be the associated blow-up building.
Then $G$ and $Y$ are quasi-isometric.
@@
Let $\alpha:H\curvearrowright\mathbb{B}_{\Gamma}$ be an action by cubical automorphisms. Let $(g_{\ell})_{\ell}$ be a blow-up datum, let $(f_{\ell})_{\ell}$ be an adapted family of bijections, and let $Y$ be the associated blow-up building.
If $\alpha$ is compatible with $(g_{\ell}),(f_{\ell})$, then there exists an action $\alpha^{\prime}:H\curvearrowright Y$ by cellular isometries such that the map $\pi:Y\to\mathbb{B}_{\Gamma}$ is $(\alpha^{\prime},\alpha)$-equivariant.
@@ 3.3 Proof of the quasi-isometry criterion
Let $\{\mathsf{v}_{1},\dots,\mathsf{v}_{n}\}$ be a (finite) set of representatives of the orbits of vertices for the action $H\curvearrowright\Gamma^{e}$. By assumption, the action $\alpha_{\mathsf{v}_{i}}$ is conjugate to an action on $\mathbb{Z}$ by uniform quasi-isometries. Therefore, by [HK18, Proposition 6.3], it is semi-conjugate to an action on $\mathbb{Z}$ by isometries. More precisely, there exist an isometric action $\gamma_{\mathsf{v}_{i}}:H_{\mathsf{v}_{i}}\curvearrowright\mathbb{Z}$, an $(\alpha_{\mathsf{v}_{i}},\gamma_{\mathsf{v}_{i}})$-equivariant surjection $g_{i}:Z_{\mathsf{v}_{i}}\to\mathbb{Z}$, and $C_{i}>0$ satisfying $|g^{-1}_{i}(n)|\leq C_{i}$ for every $n\in\mathbb{Z}$.
For every standard line $\ell$, we now define a map $g_{\ell}:\ell\to\mathbb{Z}$, following the construction in [HK18, Section 5.6]. First, if $\ell$ is a $\mathsf{v}_{i}$-line for some $i\in\{1,\dots,n\}$, we let $p_{i}:P_{\mathsf{v}_{i}}\to Z_{\mathsf{v}_{i}}$ be the projection map, and let $g_{\ell}=(g_{i}\circ p_{i})_{|\ell}$. In general $\ell$ is a $\mathsf{w}$-line for some $\mathsf{w}\in V\Gamma^{e}$. For every $\mathsf{w}\in V\Gamma^{e}$, we choose an element $h_{\mathsf{w}}\in H$ such that $h_{\mathsf{w}}\mathsf{w}=\mathsf{v}_{i}$ for some $i\in\{1,\dots,n\}$. Now if $\ell$ is a $\mathsf{w}$-line, then $h_{\mathsf{w}}(\ell)$ is a $\mathsf{v}_{i}$-line, and we let $g_{\ell}=g_{h_{\mathsf{w}}(\ell)}\circ h_{\mathsf{w}}$. As the maps $g_{i}$ are finite-to-one, the family $(g_{\ell})_{\ell}$ is a uniformly locally finite blow-up datum.
Let $(\beta_{\mathsf{v}})_{\mathsf{v}\in V\Gamma^{e}}$ be the family of branched lines associated to the blow-up datum $(g_{\ell})_{\ell}$, and let $(f_{\ell}:\ell\to t(\beta_{\mathsf{v}}))_{\ell}$ be an adapted family of bijections. For every $i\in\{1,\dots,n\}$, the equivariance property of the map $g_{i}$ gives an isometric action $\alpha^{\prime}_{\mathsf{v}_{i}}:H_{\mathsf{v}_{i}}\curvearrowright\beta_{\mathsf{v}_{i}}$. More generally, for every $\mathsf{w}\in V\Gamma^{e}$, there is an isometric action $\gamma_{\mathsf{w}}:H_{\mathsf{w}}\curvearrowright\mathbb{Z}$ given by $\gamma_{\mathsf{w}}(h)(z)=\gamma_{\mathsf{v}_{i}}(h_{\mathsf{w}}hh_{\mathsf{w}}^{-1})(z)$ (where $i$ is such that $h_{\mathsf{w}}\mathsf{w}=\mathsf{v}_{i}$). Then for every $\mathsf{w}$-line $\ell$, the map $g_{\ell}:\ell\to\mathbb{Z}$ constructed above is $(\alpha_{\mathsf{w},\ell},\gamma_{\mathsf{w}})$-equivariant. So it yields an isometric action $\alpha^{\prime}_{\mathsf{w}}:H_{\mathsf{w}}\curvearrowright\beta_{\mathsf{w}}$, such that the map $f_{\ell}:\ell\to t(\beta_{\mathsf{w}})$ is $(\alpha_{\mathsf{w},\ell},\alpha^{\prime}_{\mathsf{w}})$-equivariant. One also checks that the second compatibility condition in Definition 3.7 is satisfied.
Let $Y$ be the blow-up building associated to $(g_{\ell})_{\ell},(f_{\ell})_{\ell}$. By Lemma 3.6, the space $Y$ is quasi-isometric to $G$. By Lemma 3.8, there is an action $\alpha^{\prime}:H\curvearrowright Y$ by cellular isometries, such that the map $\pi:Y\to\mathbb{B}_{\Gamma}$ is $(\alpha^{\prime},\alpha)$-equivariant.
@@
We now prove that the action $\alpha^{\prime}$ is proper and cocompact. This will imply that $H$ is finitely generated and quasi-isometric to $G$, as desired. By definition of $\pi$, every rank 0 vertex of $\mathbb{B}_{\Gamma}$ has a unique preimage under $\pi$. As the map $\pi:Y\to\mathbb{B}_{\Gamma}$ is $(\alpha^{\prime},\alpha)$-equivariant, it follows that the action $\alpha^{\prime}$ has finitely many orbits of rank 0 vertices. Note that every vertex $y$ of $Y$ of rank at least $1$ is adjacent to at least one vertex of lower rank (this follows by considering the standard flat containing $y$ and the associated standard branched flat). Thus there exists $C>0$ such that any vertex in $Y$ is at most distance $C$ from a rank 0 vertex. As $Y$ is uniformly locally finite, it follows that $\alpha^{\prime}$ has finitely many orbits of vertices. Using again that $Y$ is uniformly locally finite, this is enough to ensure that there are only finitely many orbits of cells, so $\alpha^{\prime}$ is cocompact. As $H$ acts cocompactly on a uniformly locally finite complex, to show that the action is proper, it suffices to show that the stabilizer of each vertex is finite. The case of rank 0 vertices follows from the first assumption of the theorem, the equivariance of $\pi$, and the fact that $\pi$ is a bijection between rank 0 vertices of $Y$ and $\mathbb{B}_{\Gamma}$. The equivariance of $\pi$ ensures that the $H$-action on $Y$ preserves the rank of vertices. Therefore, the stabilizer of each vertex $y\in Y$ of rank at least $1$ permutes the non-empty finite set of vertices of lower rank which are adjacent to $y$. By induction on the rank, we thus deduce that the stabilizer of every vertex is finite.
@@ 4.2 Reduction of couplings
In this section, we review work of Kida [Kid10, Kid11] and Bader–Furman–Sauer [BFS13]. Let $L$ be a Polish group, and let $\mathsf{G}$ be a countable subgroup of $L$. Then $L$ is equipped with an action of $\mathsf{G}\times\mathsf{G}$ by left-right multiplication, namely $(g_{1},g_{2})\cdot\ell=g_{1}\ell g_{2}^{-1}$. Throughout the paper, we will use the following definition.
Definition 4.1 (Strongly ICC).
Let $L$ be a Polish group, and let $\mathsf{G}$ be a countable subgroup of $L$. The inclusion $\mathsf{G}\subseteq L$ is strongly ICC if the Dirac mass at $\mathrm{id}$ is the unique Borel probability measure on $L$ which is invariant under the conjugation by every element of $\mathsf{G}$.
@@
Let $\mathsf{H}$ be a countable group that is measure equivalent to $\mathsf{G}$, and let $\Omega$ be a measure equivalence coupling between $\mathsf{G}$ and $\mathsf{H}$.
Then there exist a homomorphism $\iota:\mathsf{H}\to L$ with finite kernel, and a measurable almost $(\mathsf{G}\times\mathsf{H})$-equivariant map $\theta:\Omega\to L$, i.e. for a.e. $\omega\in\Omega$, and any $(g,h)\in\mathsf{G}\times\mathsf{H}$, one has $\theta((g,h)\cdot\omega)=g\theta(\omega)\iota(h)^{-1}$.
@@
[…]
@@ 4.3 Restricting couplings to stabilizers
Let $L$ be a Polish group and $\mathsf{G}$ be a countable subgroup of $L$. Let $L^{\prime}\subseteq L$ be a Borel subgroup such that $\mathsf{G}\cdot L^{\prime}=L$.
Let $\mathsf{H}$ be a countable group that is measure equivalent to $\mathsf{G}$, let $(\Omega,\mu)$ be a measure equivalence coupling between $\mathsf{G}$ and $\mathsf{H}$, and assume we are given a homomorphism $\iota:\mathsf{H}\to L$ and a measurable almost $(\mathsf{G}\times\mathsf{H})$-equivariant map $\theta:\Omega\to L$, where the $(\mathsf{G}\times\mathsf{H})$-action on $L$ is via $(g,h)\cdot\ell=g\ell\iota(h)^{-1}$.
Then the groups $\mathsf{G}^{\prime}=\mathsf{G}\cap L^{\prime}$ and $\mathsf{H}^{\prime}=\iota^{-1}(L^{\prime})$ are measure equivalent. More precisely $\Omega^{\prime}=\theta^{-1}(L^{\prime})$ is a measure equivalence coupling between $\mathsf{G}^{\prime}$ and $\mathsf{H}^{\prime}$. In addition, for every subgroup $\mathsf{K}$ of either $\mathsf{G}$ or $\mathsf{H}$, every Borel fundamental domain for the action of $\mathsf{K}\cap\mathsf{G}^{\prime}$ (or $\mathsf{K}\cap\mathsf{H}^{\prime}$) on $\Omega^{\prime}$ is contained in a Borel fundamental domain for the action of $\mathsf{K}$ on $\Omega$.
@@
Let $\mathsf{K}\subseteq\mathsf{H}$ be a subgroup, let $\mathsf{K}^{\prime}=\mathsf{K}\cap\mathsf{H}^{\prime}$, and let $Y^{\prime}$ be a Borel fundamental domain for the action of $\mathsf{K}^{\prime}$ on $\Omega^{\prime}$: this exists because the $\mathsf{H}$-action on $\Omega$ has one. We claim that for any $k\in\mathsf{K}\setminus\{1\}$, one has $\mu(kY^{\prime}\cap Y^{\prime})=0$. Indeed, for a.e. $x,y\in Y^{\prime}$, if $y=kx$ for some $k\in\mathsf{K}$, then $\theta(y)=\theta(x)\iota(k)^{-1}$. As $\theta(x),\theta(y)\in L^{\prime}$, we have $k\in\mathsf{K}^{\prime}$. The claim thus follows from the fact that $Y^{\prime}$ is a fundamental domain for the action of $\mathsf{K}^{\prime}$ on $\Omega^{\prime}$. The same argument applies to subgroups of $\mathsf{G}$.
When $\mathsf{K}=\mathsf{H}$, the above claim implies in particular that every Borel fundamental domain for the action of $\mathsf{H}^{\prime}$ on $\Omega^{\prime}$ has finite measure. Likewise, any Borel fundamental domain for the action of $\mathsf{G}^{\prime}$ on $\Omega^{\prime}$ has finite measure.
@@
Let $K$ be a polyhedral complex with countably many cells. Let $L$ be a Polish group acting on $K$ through a measurable homomorphism $L\to\mathrm{Aut}(K)$. Let $\mathsf{G}$ be a countable subgroup of $L$, and assume that the actions of $\mathsf{G}$ and $L$ on $VK$ have the same orbits.
Let $\mathsf{H}$ be a countable group that is measure equivalent to $\mathsf{G}$, let $(\Omega,\mu)$ be a measure equivalence coupling between $\mathsf{G}$ and $\mathsf{H}$, and assume we are given a homomorphism $\iota:\mathsf{H}\to L$ and a measurable almost $(\mathsf{G}\times\mathsf{H})$-equivariant map $\theta:\Omega\to L$.
Then for every $v\in VK$, the groups $\mathsf{G}_{v}=\mathrm{Stab}_{\mathsf{G}}(v)$ and $\mathsf{H}_{v}=\iota^{-1}(\mathrm{Stab}_{L}(v))$ are measure equivalent. More precisely $\Omega_{v}=\theta^{-1}(\mathrm{Stab}_{L}(v))$ is a measure equivalence coupling between $\mathsf{G}_{v}$ and $\mathsf{H}_{v}$, and every Borel fundamental domain for the action of $\mathsf{G}_{v}$ (resp. $\mathsf{H}_{v}$) on $\Omega_{v}$ is contained in a Borel fundamental domain for the action of $\mathsf{G}$ (resp. $\mathsf{H}$) on $\Omega$.
@@
Given a countable set $\mathbb{D}$, we denote by $\mathrm{Bij}(\mathbb{D})$ the group of all bijections of $\mathbb{D}$; we equip it with the topology of pointwise convergence, for which it is a Polish group. Recall from the introduction that a group $\mathsf{H}$ has bounded finite subgroups if there is a bound on the cardinality of a finite subgroup of $\mathsf{H}$.
@@
Let $\mathsf{H}$ be a countable group with bounded finite subgroups that is measure equivalent to $\mathsf{G}$, and let $\Omega$ be a measure equivalence coupling between $\mathsf{G}$ and $\mathsf{H}$. Assume that we are given a homomorphism $\iota:\mathsf{H}\to L^{0}$ and a measurable almost $(\mathsf{G}\times\mathsf{H})$-equivariant map $\theta:\Omega\to L^{0}$.